OpenStudy (dtan5457):

Given f(x)=4x^2-x and g(x)=2x^2 find f(g(x))

Directrix (directrix):

f( x )= 4 * x^2 - x f ( g( x) ) = f( 2x^2 ) = 4 * (2x^2) ^2 - ( 2x^2) = ? Whatever f did to x; in this case, squared it, multiplied by 4 and then subtracted it, f does those same operations t o ( 2 x^2) which now plays the role of x.

OpenStudy (dtan5457):

i had that and ended up with 2x(4x^3-x), which is wrong

Directrix (directrix):

Let me see what I get.

Directrix (directrix):

I got 16 x^4 - 2x^2

Directrix (directrix):

Look at this again: 4 * (2x^2) ^2 - ( 2x^2) =

OpenStudy (dtan5457):

oh crud.. (2x^2)^2=4x^4... Dang, I put 2x^4 on my test.

Directrix (directrix):

Yikes.

OpenStudy (dtan5457):

So would 16x^4-2x^2 be the final answer, or should I factor the gcf?

Directrix (directrix):

I don't know that there is a set policy on that. I think that I would.